OpenStudy (anonymous):

Solve for "r": 4/3πr^3 = V I understand I need to get "r" by itself, but I'm confused on how to get the "4/3π" out of the way. Do I multiply it out? And if so, how?

OpenStudy (anonymous):

\[\frac{4}{3}\pi r^3=V,3V=4 \pi r^3\] \[\frac{3V}{4 \pi}=r^3\] \[r=\sqrt[3]{\frac{3V}{4 \pi}}\]